A man invested $6000 dollars and if 5% is in simple intrest and the rest is in 7% interest. if the anunal income on the interest is $372 how much money was put in each one?
Answer by orca(409)

Let x be the amount invested at 5%, the the amount invested at 7% is 6000 - x.
The interest earned from the 5% investment is x*5%.
The interest earned from the 7% investment is (6000 - x)*7%.
Their total is 372, we have:
x*5% + (6000 - x)*7% = 372
Solving the equation for x, we have
5x + 7(6000 - x) = 37200 (multiply both sides by 100)
5x + 42000 - 7x = 37200
-2x = -4800
x = 2400
So
the amount invested at 5% is $2400.
the amount invested at 7% is 6000 - x = 6000 - 2400 = $3600.